OpenStudy (anonymous):

Write a polynomial equation with integer coefficients that has the given roots. x=-8 x=7

TheSmartOne (thesmartone):

so the roots have to be -8 and 7 ?

OpenStudy (anonymous):

yes

TheSmartOne (thesmartone):

To find the equation it should be in the form (x-_)(x-_)=0 We have to plug in -8 and 7 into it.

OpenStudy (anonymous):

ok lets see

TheSmartOne (thesmartone):

And then we multiply it out to get the equation.

OpenStudy (anonymous):

so it s(x-8)(x-7)=0 ?

TheSmartOne (thesmartone):

So (x-(-8))(x-7)=0

TheSmartOne (thesmartone):

x-(-8)=x+8

TheSmartOne (thesmartone):

@beautii

OpenStudy (anonymous):

my apology , I am a bit confused

TheSmartOne (thesmartone):

Because x+8=0 then x=-8

OpenStudy (anonymous):

the polynomial equation is x-(-8)=x+8 ?

OpenStudy (anonymous):

x+8=0 x=-8 what about x=7 ?

TheSmartOne (thesmartone):

so then to make x=7 just subtract 7 we get x-7=0

TheSmartOne (thesmartone):

so now we multiply (x-7)(x+8)=0

TheSmartOne (thesmartone):

Also \(\Huge\cal\color{red}{Welcome~to~Open~Study!!}\) and since you are new here feel free to look at this. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

@beautii

OpenStudy (anonymous):

thanks so much

TheSmartOne (thesmartone):

No problem